# Optimal Online Change Detection via Random Fourier Features

### v1

@@ A.4.5 Proof of Lemma 1
$\leq 2\exp\left(-\frac{1}{2}n\varepsilon^{2}\left[\sigma^{2}+2\sqrt{2}\varepsilon\right]^{-1}\right)$
where the final line holds due to Jensen’s inequality and the fact that
$\mathbb{E}\left[\tilde{\sigma}^{2}\left(\omega_{1},\dots,\omega_{r}\right)\right]$
[…]
=$=\left\{1+\left|I_{i^{\prime}}\right|\left\lfloor\frac{m}{n}\right\rfloor^{-2}\right\}\mathbb{E}_{X\sim\mathbb{P}}\left[K\left(X,X\right)\right]$
@@
$\quad-\left\{2\left|I_{i^{\prime}}\right|\left\lfloor\frac{m}{n}\right\rfloor^{-1}-\left|I_{i^{\prime}}\right|\left(\left|I_{i^{\prime}}\right|-1\right)\left\lfloor\frac{m}{n}\right\rfloor^{-2}\right\}\mathbb{E}_{X,Y\sim\mathbb{P}}\left[K\left(X,Y\right)\right]$
$\leq 2\mathbb{E}_{X\sim\mathbb{P}}\left[K\left(X,X\right)\right]-\mathbb{E}_{X,Y\sim\mathbb{P}}\left[K\left(X,Y\right)\right]=\sigma^{2}.$
This proves the desired result. ∎
@@ A.5 Auxiliary results
In this section, we collect a few auxiliary results. Besides establishing useful bounds on real numbers in Lemma 2 and Lemma 3, we show that the self-bounding property of RFF-MMD (Lemma 4) leads to its exponential concentration (Lemma 5). The latter is one of the key ingredients for deriving our threshold sequences elaborated Section 4.3.
@@
The RFF-MMD as defined in (9) between two empirical measures composed respectively of $n$ and $m$ sample points is a function mapping from $(\mathbb{R}^{d})^{m+n}\to\mathbb{R}$. This function has the self bounding property with constants
@@
$\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{K}\right]=\sup_{f\in\mathcal{H}_{K}:\left\|f\right\|_{\mathcal{H}_{K}}\leq 1}\left(\mathbb{E}_{X\sim\mathbb{P}}\left[f\left(X\right)\right]-\mathbb{E}_{Y\sim\mathbb{Q}}\left[f\left(Y\right)\right]\right)$
@@
$\leq\sup_{\begin{subarray}{c}\mathbf{x},\tilde{\mathbf{x}}\\ f\in\hat{\mathcal{G}}\end{subarray}}\frac{1}{n}\left|f(\mathbf{x})-f(\tilde{\mathbf{x}})\right|=\sup_{\mathbf{x},\tilde{\mathbf{x}}}\frac{1}{n}\left\|\hat{K}\left(\cdot,\mathbf{x}\right)-\hat{K}\left(\cdot,\tilde{\mathbf{x}}\right)\right\|_{\hat{\mathcal{H}}}$
@@
$\hat{K}\left(\cdot,\mathbf{x}\right),\hat{K}\left(\cdot,\mathbf{x}\right)\right>_{\mathcal{H}_{\hat{K}}}^{2}=\cos(0)=1.$
@@
$=\mathbb{E}\left[\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\hat{K}\left(X_{i},X_{j}\right)+\frac{1}{m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\hat{K}\left(Y_{i},Y_{j}\right)\right.\right.$
@@
($\leq\Bigg{(}\left\{\frac{1}{n}+\frac{1}{m}\right\}\underbrace{\mathbb{E}\left[\hat{K}\left(X,X\right)\mid\omega_{1},\dots,\omega_{r}\right]}_{=1}$
+$\hskip 120.00018pt+\underbrace{\left\{\frac{n-1}{n}+\frac{m-1}{m}-2\right\}}_{=-\left(\frac{1}{n}+\frac{1}{m}\right)}\mathbb{E}\left[\hat{K}\left(X,Y\right)\mid\omega_{1},\dots,\omega_{r}\right]\Bigg{)}^{\frac{1}{2}}$
=($=\Bigg{(}\left\{\frac{1}{n}+\frac{1}{m}\right\}\Bigg{(}1-\underbrace{\mathbb{E}\left[\hat{K}\left(X,Y\right)\mid\omega_{1},\dots,\omega_{r}\right]}_{\geq-1}\Bigg{)}\Bigg{)}^{\frac{1}{2}}$
@@ A.6 External statements
In this section, we collect the external statements that we use, to ensure self-completeness. Theorem 5 recalls McDiarmid’s inequality from Boucheron et al. [5, Section 6.1], which is also known as bounded differences inequality [34]. Theorem 6 is about the concentration of random Fourier features and part of the proof of Sriperumbudur and Szabó [50, Theorem 1]. We recall the concentration result Yurinsky [60, Theorem 3.3.4] on random variables taking values in a separable Hilbert space in Theorem 7.
@@
Then, if $X_{1},\dots,X_{n}$ is a sequence of identically distributed random variables and (38) holds, putting $Z=f\left(X_{1},\dots,X_{n}\right)$ and $\nu=\frac{1}{4}\sum_{i=1}^{n}c_{i}^{2}$ for any $t>0$, it holds that
@@
Let $\hat{K}$ be defined as in (8). Let $\mathcal{X}$ a proper subset of $\mathbb{R}^{d}$ and denote by $|\mathcal{X}|$ its Lebesgue measure. For any $t>0$, it holds that
@@
where $\sigma^{2}=\int\left\|\omega\right\|_{2}^{2}\mathrm{d}\Lambda\left(\omega\right)$ and
@@
[…]

### v2

@@ A.4.6 Proof of Theorem 5
$A_{2}=\left\{\sup_{\mathbf{x},\mathbf{y}\in\mathcal{X}}\left|\hat{K}\left(\mathbf{x},\mathbf{y}\right)-K\left(\mathbf{x},\mathbf{y}\right)\right|\leq\frac{h\left(d,\left|\mathcal{X}\right|,\sigma\right)+\sqrt{2\log(2/\alpha)}}{\sqrt{r}}\right\},$ (39)
on which the random Fourier feature kernel $\hat{K}$ is close to $K$ in sup norm sense. Note that by Theorem 7 the event (39) holds with probability at least $1-\alpha/2$. Introduce also the sequence of events
@@
$B_{k}=\left\{\eta_{k}<\hat{\eta}_{k}\leq\eta_{k}+C_{3}\frac{\log\left(T/\alpha\right)}{\left(\operatorname{MMD}_{K}\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\right)^{2}}\right\},\quad\text{for}\quad k=1,\dots M,$
such that on the event $B_{k}$ the $k$-th change point is detected and localized at the required rate. If $\hat{M}<M$ we employ the convention $\hat{\eta}_{k}=T$ for all $k>\hat{M}$. We now show that on (37) and (39) the event $B_{1}$ holds with probability at least $T/(2\alpha)$. Introduce the quantity
$n_{1}^{*}=\min\left\{n<\eta_{2}\mid n-\eta_{1}=2^{j_{1}^{*}}\text{ for }j_{1}^{*}\in\mathbb{N},\sqrt{2^{j_{1}^{*}}}\text{MMD}_{K}[\mathbb{Q}_{1},\mathbb{Q}_{2}]>\sqrt{C_{3}\log(2T/\alpha)}\right\}$
To ease the notation put $I_{1}^{*}=1:\eta_{1}$, $J_{1}^{*}=(\eta_{1}+1):n_{1}^{*}$, and $V_{1}^{*}=\frac{2^{j_{1}^{*}}(n_{1}^{*}-2^{j_{1}^{*}})}{n_{1}^{*}}$. Observe that
$\sqrt{V_{1}^{*}}\text{MMD}_{\hat{K}}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]$
$\quad\geq\sqrt{V_{1}^{*}}\text{MMD}_{K}\left[\mathbb{Q}_{1},\mathbb{Q}_{2}\right]+\sqrt{V_{1}^{*}}\left(\text{MMD}_{K}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]-\mathbb{E}\left[\text{MMD}_{K}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\right]\right)$ (40a)
$\quad\quad-2\sqrt{V_{1}^{*}}\sqrt{\sup_{\bm{x,y}}\left|\hat{K}(\bm{x,y})-K(\bm{x,y})\right|}$
$\quad\quad-\sqrt{V_{1}^{*}}\left|\mathbb{E}\left[\text{MMD}_{K}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\right]-\text{MMD}_{K}\left[\mathbb{Q}_{1},\mathbb{Q}_{2}\right]\right|$
$\quad\geq\frac{1}{2}\sqrt{V_{1}^{*}}\text{MMD}_{K}\left[\mathbb{Q}_{1},\mathbb{Q}_{2}\right]+\sqrt{V_{1}^{*}}\left(\text{MMD}_{K}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]-\mathbb{E}\left[\text{MMD}_{K}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\right]\right)-\frac{4}{\sqrt{2}}$ (40b)
Where (40a) holds by the triangle inequality and arguments similar to (21), and (40b) holds by arguments similar to (22) together with the fact that on (39) we will have that
$\sup_{\bm{x,y}}\left|\hat{K}(\bm{x,y})-K(\bm{x,y})\right|\leq\frac{1}{4}\min_{k=1,\dots,M}\left(\text{MMD}_{K}\left[\mathbb{Q}_{k},\mathbb{Q}_{k+1}\right]\right)^{2}\leq\frac{1}{4}\left(\text{MMD}_{K}\left[\mathbb{Q}_{1},\mathbb{Q}_{2}\right]\right)^{2}.$
Using the above we therefore have that
$\mathbb{P}\left(B_{1}^{c}\right)$ $\leq\mathbb{P}\left(\sqrt{V_{1}^{*}}\text{MMD}_{\hat{K}}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\leq\lambda_{T}\right)$
$\leq\mathbb{P}\left(\sqrt{V_{1}^{*}}\left(\mathbb{E}\left[\text{MMD}_{\hat{K}}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\right]-\text{MMD}_{\hat{K}}\left[X_{I_{1}^{*}},X_{J_{1}^{*}}\right]\right)\right$.
$\quad\quad\quad\quad\quad\left.>\frac{1}{2}\sqrt{V_{1}^{*}}\text{MMD}_{K}\left[\mathbb{Q}_{1},\mathbb{Q}_{2}\right]-\left(\sqrt{4}+\sqrt{6}\right)\sqrt{\log(2T/\alpha)}\right)\leq\frac{\alpha}{2T},$
where the final inequality follows from the definition of $C_{3}$, the bounded difference property of $\text{MMD}_{K}$, and Theorem 6. By the definition we must have that $2^{j_{1}^{*}}\leq\delta_{1}/2$, therefore on the event $A_{1}$ the first change point is detected and the algorithm starts at most mid way between $\eta_{1}$ and $\eta_{2}$. Identical arguments to those above therefore give that, on the events $A_{1}$, $A_{2}$, and $B_{1}$ the event $B_{2}$ holds with probability at least $1-\alpha/(2T)$ and having detected the second change point the algorithm re-starts at most mid way between $\eta_{2}$ and $\eta_{3}$. By induction and a union bound argument, on the events (37) and (39) the events $B_{1},\dots B_{M}$ hold with probability at least $1-M\alpha/(2T)\geq 1-\alpha/2$. Since (37) and (39) jointly hold with probability $1-3\alpha/2$ we are done. ∎
@@ A.5 Auxiliary results
In this section, we collect a few auxiliary results. Besides establishing useful bounds on real numbers in Lemma 2 and Lemma 3, we show that the bounded differences property of RFF-MMD (Lemma 4) leads to its exponential concentration (Lemma 5). The latter is one of the key ingredients for deriving our threshold sequences elaborated Section 4.3.
@@
The RFF-MMD as defined in (7) between two empirical measures composed respectively of $n$ and $m$ sample points is a function mapping from $(\mathbb{R}^{d})^{m+n}\to\mathbb{R}$. This function has the bounded differences property with constants
@@
$\operatorname{MMD}_{K}\left[\mathbb{P},\mathbb{Q}\right]=\sup_{f\in\mathcal{H}_{K}:\left\|f\right\|_{\mathcal{H}_{K}}\leq 1}\left(\mathbb{E}_{X\sim\mathbb{P}}\left[f\left(X\right)\right]-\mathbb{E}_{Y\sim\mathbb{Q}}\left[f\left(Y\right)\right]\right)$
@@
[…]
@@
$\left<\hat{K}\left(\cdot,\mathbf{x}\right),\hat{K}\left(\cdot,\mathbf{x}\right)\right>_{\mathcal{H}_{\hat{K}}}^{2}=\cos(0)=1.$
@@
[…]
@@
$\leq\Bigg(\left\{\frac{1}{n}+\frac{1}{m}\right\}\underbrace{\mathbb{E}\left[\hat{K}\left(X,X\right)\mid\omega_{1},\dots,\omega_{r}\right]}_{=1}$
$\hskip 120.00018pt+\underbrace{\left\{\frac{n-1}{n}+\frac{m-1}{m}-2\right\}}_{=-\left(\frac{1}{n}+\frac{1}{m}\right)}\mathbb{E}\left[\hat{K}\left(X,Y\right)\mid\omega_{1},\dots,\omega_{r}\right]\Bigg)^{\frac{1}{2}}$
$=\Bigg(\left\{\frac{1}{n}+\frac{1}{m}\right\}\Bigg(1-\underbrace{\mathbb{E}\left[\hat{K}\left(X,Y\right)\mid\omega_{1},\dots,\omega_{r}\right]}_{\geq-1}\Bigg)\Bigg)^{\frac{1}{2}}$
@@ A.6 External statements
In this section, we collect the external statements that we use, to ensure self-completeness. Theorem 6 recalls McDiarmid’s inequality from Boucheron et al. [5, Section 6.1], which is also known as bounded differences inequality [38]. Theorem 7 is about the concentration of random Fourier features and part of the proof of Sriperumbudur and Szabó [54, Theorem 1]. We recall the concentration result Yurinsky [65, Theorem 3.3.4] on random variables taking values in a separable Hilbert space in Theorem 8.
@@
Then, if $X_{1},\dots,X_{n}$ is a sequence of independently distributed random variables and (44) holds, putting $Z=f\left(X_{1},\dots,X_{n}\right)$ and $\nu=\frac{1}{4}\sum_{i=1}^{n}c_{i}^{2}$ for any $t>0$, it holds that
@@
Let $\hat{K}$ be defined as in (6). Let $\mathcal{X}$ be a proper subset of $\mathbb{R}^{d}$ and denote by $|\mathcal{X}|$ its Lebesgue measure. For any $t>0$, it holds that
@@
where $\sigma^{2}=\int_{\mathbb{R}^{d}}\left\|\omega\right\|_{2}^{2}\mathrm{d}\Lambda\left(\omega\right)$ and
@@
Let $X_{1},\dots,X_{n}$ be a sequence of zero mean independent random variables taking values in a real and separable Hilbert space $\mathcal{X}$ with inner product $\left$$\cdot$$\cdot\right$> and norm $\left\|\cdot\right\|=\sqrt{\left<\cdot,\cdot\right>}$. Write $S_{n}^{*}=\sup_{m\leq n}\left\|X_{1}+\dots+X_{m}\right\|$. If the random variables satisfy the moment condition